OpenStudy (sasogeek):

I'm beginning to understand integration... anyone mind giving me something to try out...you guys got questions of which u know the answers... i wanna try them too so let me have them and let me know if i make a mistake... :/

OpenStudy (anonymous):

integrate cos 3x with respect to x

OpenStudy (anonymous):

\[\int_0^{\infty} 2^t e^t dt\]

OpenStudy (sasogeek):

jim: 1/3 sin3x + c

OpenStudy (anonymous):

yup

OpenStudy (lalaly):

\[\int\limits{24x \ln( 6x^2) dx}\]

OpenStudy (anonymous):

want an improper??

OpenStudy (anonymous):

\[\huge \int_{1}^{2}\frac{cosh(lnt)}{t}dt\]

OpenStudy (anonymous):

try this \[\int\limits_{?}^{?} 1/x\]

OpenStudy (lalaly):

dx

OpenStudy (sasogeek):

i was trying to solve lana's question but i'm stuck at solving the natural log... need help

OpenStudy (anonymous):

get mine?

OpenStudy (lalaly):

by parts saso

OpenStudy (sasogeek):

imran i can't do ur question... the very first one... and lana: int dx = x

OpenStudy (anonymous):

anyone?

OpenStudy (lalaly):

no lol there was a dx missing in that persons integral

OpenStudy (anonymous):

\[\int\limits_{0}^{1}2x \div(x ^{2}+3) \]

OpenStudy (sasogeek):

but is the integration of dx = x? was i right?

OpenStudy (lalaly):

yup:D

OpenStudy (anonymous):

what about the intergration of 1/x?

OpenStudy (sasogeek):

ln x

OpenStudy (sasogeek):

btw though i don't know how to solve integrals with a linear function involving fractions where the variable is a part of the denominator... like jimmy's last question... i'm going to learn that later today or better still someone could explain it to me now...

OpenStudy (anonymous):

try mine. It's just algebra

OpenStudy (sasogeek):

imran i'm going to try and solve urs on a sheet of paper first before trying to type it out so you'd have to wait for a bit :)

OpenStudy (anonymous):

for jimmys you can is U sub

OpenStudy (anonymous):

so u = \[u = x^{2}+3\] Du = 2x

OpenStudy (anonymous):

i think thats it im not sure i jus got out of a 3 hour exam and am tired lol

OpenStudy (sasogeek):

imran the answer is 2

OpenStudy (sasogeek):

right?

OpenStudy (zarkon):

\[\int_0^{\infty} 2^t e^t dt=\infty\]

OpenStudy (anonymous):

\[\[2^t= e^{t *log(2)}\] \[\int_0^{\infty} e^{t *log(2)} * e^t dt\] \[\int_0^{\infty} e^{t *log(2)+t}dt \] \[\int_0^{\infty} e^{t (*log(2)+1)} \]

OpenStudy (anonymous):

I know we could have used comparison test, but this is for any limits

OpenStudy (zarkon):

\[2^t e^t \ge1 \text{ for }t\ge 0\] thus \[\int_0^{\infty} 2^t e^t dt\ge \int_0^{\infty} 1 dt=\infty\]

OpenStudy (sasogeek):

\[\int\limits_{0}^{\infty}2^{t}e ^{t}dt=\left[ ((2^{t+1})/t+1)e ^{t} \right]_{0}^{\infty}\]

OpenStudy (zarkon):

I would just write \[2^te^t\]as\[(2e)^t\]

OpenStudy (zarkon):

use \[\int a^tdt=\frac{a^t}{\ln(a)}+c\]

OpenStudy (anonymous):

I got this out of finding laplace transform of 2^t

OpenStudy (sasogeek):

omg i made a huge mistake lol

OpenStudy (sasogeek):

the 2 was a base to t, not a coefficient so i integrated so wrongly

OpenStudy (sasogeek):

facepalm

OpenStudy (sasogeek):

zarkon i haven't seen ur formula before though... there's a lot i'm yet to learn :P i didn't expect this level of questions lol but it was worth seeing what the future beholds for me